OpenStudy (anonymous):

What are the dimensions of a rectangle that has a perimeter of 148 units and an area of 144 square units? a. 3 by 48 c. 37 by 37 b. 2 by 72 d. 4 by 36

jhonyy9 (jhonyy9):

perimeter equal 2w+2l=148 area equal w*l=144 2w+2l=148 2(w+l=148 divide both sides by 2 w+l=74 ---- w=74-l this substitute inside second equation in place of wand will get (74-l)*l = 144 74l-i^2 =144 and have got an quadratic in l what need to solve it l^2 -74l +144 =0 factorized this is equal (l-2)(l-72) =0 l-2=0 l=2 l-72=0 l=72 this mean that the length of rectangle is 72 and the width 2 units hope so much that is understandably easy

jhonyy9 (jhonyy9):

b. will be right from choices